You own the following portfolio of stocks. What is the portfolio weight of Stock C?

Stock Number of Shares Price per Share
A 120 $ 27
B 700 $ 23
C 400 $ 47
D 200 $ 46

Multiple Choice

  • .0684

  • .3401

  • .1943

  • .6587

  • .3971
Answer #1
Stock Total value=(Number of shares*Price per share)
A (120*27)=3240
B (700*23)=16100
C (400*47)=18800
D (200*46)=9200
Total=47340

Hence portfolio weight of Stock C=18800/47340

=0.3971(Approx).